OpenStudy (anonymous):

Function Table Help I will put the Input values first, then the output. X: 1, 2, 3, 4, 5 Y: 2, -3, -8, -13 X: 1, 2, 3, 4, 5 Y: 3, -1, -5, -9 I am really stuck on finding the function for these two.

OpenStudy (anonymous):

the first one is the relation is X is +1 ex: 1+2=3 and Y is -5 ex: 2-5= -3 the second one is X is +1 and Y is -4 I hope this helped!

OpenStudy (anonymous):

Im not sure this is correct. I was not clear on this but you need to find the relationship between the Y and the X.

OpenStudy (anonymous):

I think that that is the first step to finding a relationship between the two i guess you it can be a linear function because it is not a quadratic function and is not an exponential function

OpenStudy (debbieg):

In the first one, each increase in x results in a decrease of 5 for y. So y=-5x+b To find b, use the first pair, (1,2): 2=-5(1)+b b=7 So y=-5x+7 Now do the same thing for the 2nd one - look at the changes in the y's as x changes by 1. It is again linear, since the change for a 1 unit increase in x is the same.

OpenStudy (anonymous):

This is what I was looking for! Thank you so much! Nice try though Rose.

OpenStudy (anonymous):

no problem sorry i couldn't help :)